A cyclist starts from rest and accelerates at 2m%2Fs square for 20 seconds. He then travels at a constant speed for 1 minute and then finally decelerates at 3m%2Fs square. Find the maximum speed
1
Expert's answer
2018-07-18T15:38:08-0400
Since the speed of the cyclist decreases during the deceleration, the maximum speed he has after the acceleration (after first 20 seconds and during one minute after).
The initial speed v_i, final speed v_f, acceleration a and the time of acceleration t are related by the following expression
v_f-v_i=at
Since v_i=0 m/s, a=2 m/s^2 and t=20 s, we obtain
v_f=v_i+at=0 m/s+2 m/s^2 ∙20 s=40 m/s